Kumar spends all his monthly income of $20 on two goods, rice (x) and cooking oil (y). The price of oil is $1 per liter. Rice can be purchased at a government-run store and also on the free market. The price of rice is $1 per kilogram at the government store but he can only buy up to 10 kgs. In the free market, the price of rice is $2 per kg. Draw Kumar’s budget constraint assuming that the goods are divisible.


1
Expert's answer
2022-01-17T12:27:01-0500

1)"Y=\frac{20}{1}-\frac{1}{2}x"


Y=20-0.5x


2)"Y=\frac{20}{1}-\frac{1}{1}x"


Y=20-x